# Optimal CO2 storage management considering safety constraints in multi-stakeholder multi-site CCS projects: a game theoretic perspective

### v1

@@ Abstract
In this work, we propose a paradigm based on Markov games to quantitatively investigate how different coalition structures affect the goals of stakeholders. We frame this multi-stakeholder multi-site problem as a multi-agent reinforcement learning problem with safety constraints. Our approach enables agents to learn optimal strategies while compliant with safety regulations. We present an example where multiple operators are injecting $CO{{}_{2}}$ into their respective project areas in a geologically connected basin. To address the high computational cost of repeated simulations of high fidelity models, a previously developed surrogate model based on the Embed-to-Control (E2C) framework is employed. Our results demonstrate the effectiveness of proposed framework in addressing optimal management of $CO{{}_{2}}$ storage when multiple stakeholders with various objectives and goals are involved.
@@
[…]
[label1]organization=Bureau of Economic Geology, Jackson School of Geosciences,
University of Texas at Austin, addressline=Austin, state=TX, postcode=78758, country=USA
@@ 1.1 Background
Geological carbon storage (GCS) has emerged as a viable technology in global efforts to mitigate climate change by permanently sequestering carbon dioxide (CO2) in deep subsurface formations. As part of the broader carbon capture and storage (CCS) strategy, GCS plays a critical role in reducing anthropogenic emissions from industrial sources while enabling continued use of fossil-based infrastructure during the energy transition (Celia et al., 2015; Hosseini et al., 2024). Modern GCS projects are increasingly being implemented at the basin scale, where geological formations span large areas and support long-term operations and large-volume storage (Huang et al., 2014; Birkholzer and Zhou, 2009; Wijaya et al., 2024). These large-scale deployments often involve multiple injection sites, complex monitoring and regulatory requirements, and diverse stakeholder participation. In this context, multi-stakeholder CCS initiatives—where operators, landowners, regulators, and investors interact within geologically connected sites—are becoming increasingly common.
@@
In a multi-stakeholder CCS setting, each stakeholder pursues distinct interests and goals (Bentham et al., 2014). For example, operators typically manage one or more injection/extraction wells with objectives such as maximizing CO2 storage or financial returns. Landowners focus on maximizing land use value and royalty income, while regulators prioritize long-term safety, pressure containment, and environmental compliance, e.g. Class VI well regulations (U.S. Environmental Protection Agency, 2013; Leng et al., 2024). Although only operators physically interact with the reservoir, interests of all other stakeholders also shape the operational landscape. For instance, regulators may impose safety constraints that limit injection volumes and rates, whereas landowners and operators may both seek to exploit the available pressure space for maximum economic return. Moreover, the interconnected nature of subsurface dynamics means that one operator’s injection strategy can alter reservoir pressure and plume behavior, influencing the economic and operational outcomes of other operators.
@@
These inter-dependencies pose significant challenges to effective basin-wide coordination. Without a robust decision-making framework that integrates individual or collective objectives with system-wide regulatory constraints, stakeholders may risk adopting locally optimal strategies that undermine collective performance, leading to underutilized reservoir capacity, reduced economic returns, or even regulatory noncompliance. To this end, coordinated multi-agent optimization is essential to balance return, safety, and fairness in basin-scale GCS deployment.
1.2 Challenges of Multi-agent Optimization in GCS
@@
The challenges of optimal management of basin-scale geological carbon storage are multifold, especially when multiple stakeholders with respective goals and interests are involved. On the one hand, high fidelity models that predict the pressure evolution and plume migration are needed to ensure meeting safety constraints. Optimization of this system requires multiple forward evaluations of these models, which becomes infeasible when the computational resources are limited. A remedy to this is employing machine learning-based proxy models to amortize the computation costs (Wang et al., 2024; Gurwicz et al., 2025). On the other hand, the literature on multi-agent optimization for GCS or broader subsurface communities is scarce. Existing literature either only explored multi-objective optimization of one single stakeholder on various CO2 storage projects (You et al., 2020; Park et al., 2021; Liu et al., 2025) or constrained co-optimization of one single stakeholder (Zou and Durlofsky, 2023; Nguyen et al., 2024; Tang and Durlofsky, 2025). Recently, (Pettersson et al., 2024, 2025) proposed a multi-objective optimization framework that enables coordinated CO2 injection strategies among multiple agents in basin-scale carbon storage projects.While this approach facilitates joint optimization, it assumes the presence of a single centralized decision maker who assigns operational strategies to all agents. In doing so, the solution selection may inherently prioritize certain objectives over others—for instance, granting more weight to one agent’s goal. This centralized assumption overlooks the reality that, in practice, each agent seeks to maximize its own returns, therefore limiting the framework’s applicability to competitive multi-operator environments, where decision-making is inherently decentralized and driven by individual objectives.
@@ 1.3 Constrained Markov Game (CMG) and Safe Multi-agent Reinforcement Learning (safe MARL)
In addition to the traditional optimization framework, which based on gradient-based or evolutionary algorithms, a framework referred to as constrained Markov game (CMG) (Altman, 2021) is suitable for constrained optimization and offers real-time optimized solutions when multiple actors are involved. A constrained Markov game extends the Markov game framework by adding safety constraints that all agents must comply with while optimizing their own objectives. In a CMG, the system is represented as a sequential decision-making process defined by states, actions, transitions, and rewards, along with explicit safety constraints that limit feasible actions. Safe multi-agent reinforcement learning provides practical methods to learn policies in such settings, ensuring agents maximize rewards while avoiding constraint violations. The detailed formulation of this approach is presented in section 3 and section 4.
@@ 1.4 Contributions and Paper Organization
The contributions of this study are twofold. First, we develop a safe multi-agent reinforcement learning framework tailored for multi-stakeholder, multi-site CO2 storage problems when each stakeholder has its own goals and interests. The framework accommodates these divergent interests while ensuring safe operations, and its performance is validated against traditional multi-objective optimization approaches. Second, a reward and penalty model is meticulously designed for CO2 storage projects, and various coalition structures based on shared rewards and penalties have been investigated to evaluate their impacts on stakeholder objectives and overall system performance.
The paper is organized as follows. Section 2 reviews the subsurface dynamics of GCS projects, including key objectives, operational constraints, and potential coalition strategies among stakeholders. Section 3 introduces fundamental concepts in reinforcement learning and Markov games, establishing the theoretical basis for our approach. Section 4 presents the proposed methodology and workflow. Section 5 reports the case study results. Finally, the paper concludes with a discussion of key findings and a summary of contributions and directions for future work.
@@ 2.1 Overview of Basin-Scale GCS System
Geological CO2 storage typically consists of studies and analyses across various scales (Guo et al., 2022; Ni et al., 2021) and involves complex coupled multi-physical processes (Nordbotten and Celia, 2011; Jiang, 2011). Deep saline aquifers have abundant storage potential and are regarded as reliable and feasible storage sites. When CO2 is compressed and supercritical CO2 is being injected into deep saline aquifers, a compositional multi-phase system is formed. In our study, we consider a compositional porous flow process where the governing equation of the system can be described as:
@@
After spatial and temporal discretizations, the above governing equations that describe the compositional flow in porous media can be reformulated as state-space form:
@@
where $x$ denotes the system states, including pressure and CO2 mole fraction, where the dimension may be too large-tens of thousands to millions, which is infeasible for off-policy RL training.
To address the computational burden of high-fidelity reservoir simulations, we leverage a surrogate reduced-order model based on the Embed-to-Control (E2C) architecture (Watter et al., 2015; Jin et al., 2020; Coutinho et al., 2021; Chen et al., 2024a, 2025b), which captures essential system dynamics while significantly reducing simulation time. Employing the machine learning-based reduced-order model, equation 3 can be approximated as:
@@
where $Enc$, DecDec, $f_{z}$ and $g_{z}$ are neural networks parameterized by $\varphi$, $\theta$, $\omega_{1}$ and $\omega_{2}$, respectively. These neural networks are trained to minimize the reconstruction, regularization and observation data losses. Prior studies have shown that this proxy model maintains high robustness while achieving desirable accuracy compared to high-fidelity simulations. For further implementation and validation details, readers are referred to (Chen et al., 2024b).
@@ 2.2 Stakeholders Objectives and Operational Conflicts
Basin-scale carbon capture and storage (CCS) projects involve a diverse array of stakeholders whose interests, responsibilities, and operational controls are inherently interdependent. Figure 1 illustrates the key actors and their roles in a typical CCS project. At the center of the system are operators, who manage CO2 injection wells (possibly also some brine extraction wells) and directly interface with the subsurface reservoir. Their primary objectives typically include maximizing CO2 storage volumes and financial returns through strategic control of injection rates and well operations.
@@
These stakeholders operate within a shared geologic domain and sometimes have misaligned goals. For instance, while both operators and landowners may favor maximizing pressure space for economic benefit, regulators may impose strict injection limits to prevent caprock fracturing or fluid migration beyond permitted boundaries. Moreover, one operator’s injection strategy can affect reservoir pressure distributions and CO2 plume migration patterns, potentially undermining neighboring operators’ performance or violating shared safety constraints.
@@ 2.3 Mathematical Formulation of Objectives
As mentioned earlier, the landlord and operators usually share the common interests, which is to maximize the financial return. Regulator’s interests can translate into safety constraints for the operators. To this end, the following objectives and constraints are proposed.
The financial return is represented as present value (PV), for each agent i, the present value/immediate reward can be defined as:
$\text{PV}_{i}(u)=\Bigg{(}$ $\sum_{j=1}^{N_{\text{inj}}^{i}}(R_{\text{credit}}-R_{\text{op}})*q_{t,j}^{i}(u)$ (5)
@@
$\text{NPV}_{i}=\sum_{t=1}^{N_{t}}(\gamma)^{t}\cdot\text{PV}_{i}$ (6)
@@
$\text{p}_{grid,i}\leq\alpha_{i}*p_{frac,i}$ (7)
@@ 2.3.1 Multi-objective optimization formulation
Intuitively, optimal CO2 storage under multi-stakeholder settings can be formulated as a constrained multi-objective optimization problem, where the objective is to maximize all companies’ NPVs while adhering to the pressure constraints. The constrained multi-objective optimization problem can be formulated as:
$\max_{u}$ $\left\{\text{NPV}_{1}(u),\,\text{NPV}_{2}(u),\,\ldots,\,\text{NPV}_{n}(u)\right\}$ (8)
@@
$p_{grid,i}(u)\leq\alpha*p_{frac,i},i=1,2,.,n$
@@
where $NPV_{i}$ indicates the NPV of agent $i$, and the decisions $u$ are bounded by operational limits, pressure constraints, and the governing system transition dynamics.
@@
Figure 2: Sketch plot of MOO with two decision variable and two objectives. The gray areas represents the feasible decision and solution space. The blue dots in the objective space represent non-dominated solutions forming the Pareto front: no point on this frontier can be improved in one objective (agent’s reward) without hurting another. The black square indicates a Nash Equilibrium: a point where neither agent can improve their individual reward by unilaterally deviating. Adapted from (Pereira et al., 2022)
@@
Traditionally, MOO problems are addressed using evolutionary algorithms to obtain a Pareto front 2, which provides a spectrum of optimal trade-offs for decision-making. One widely used method is the Non-dominated Sorting Genetic Algorithm II (NSGA-II) (Deb and Jain, 2013), implemented here using the PyMOO framework (Blank and Deb, 2020). In this approach, the algorithm iteratively evolves a population of solutions to converge toward the Pareto-optimal set. Each point on the Pareto front represents a trade-off where no objective can be improved without degrading another. This front can be used for post-optimization solution selection, typically choosing a “knee point” or preference-driven solution.
@@
Despite its effectiveness, the conventional MOO framework may not fully capture the operational realities of CCS projects, as it assumes a single centralized decision maker with complete knowledge of basin-scale reservoir conditions, inter-well interactions, and the operational strategies of all stakeholders. In practice, CCS operations often involve multiple independent operators, each with limited information and potentially competing interests, making such centralized control unrealistic. Furthermore, while CMOO yields a Pareto front, the resulting solutions often reflect inherent preferences toward certain objectives, which may advantage specific stakeholders. In real-world field operations, post-hoc optimization is rarely feasible due to regulatory, contractual, and operational constraints. These limitations reduce the representativeness of traditional CMOO in modeling actual multi-operator CCS decision-making environments.
2.3.2 Multi-agent reinforcement learning formulation

### v2

# Optimal CO2 storage management considering safety constraints in multi-stakeholder multi-site CCS projects: a Markov game perspective

@@ Abstract
In this work, we propose a paradigm based on Markov games to quantitatively investigate how different coalition structures affect the goals of stakeholders. We frame this multi-stakeholder multi-site problem as a multi-agent reinforcement learning problem with safety constraints. Our approach enables agents to learn optimal strategies while compliant with safety regulations. We present an example where multiple operators are injecting CO2 into their respective project areas in a geologically connected basin. To address the high computational cost of repeated simulations of high fidelity models, a previously developed surrogate model based on the Embed-to-Control (E2C) framework is employed. Our results demonstrate the effectiveness of proposed framework in addressing optimal management of CO2 storage when multiple stakeholders with various objectives and goals are involved.
@@
journal: International Journal of Greenhouse Gas Control
@@ 1.1 Background
Geological carbon storage (GCS) has emerged as a viable technology in global efforts to mitigate climate change by permanently sequestering carbon dioxide (CO2) in deep subsurface formations. As part of the broader carbon capture and storage (CCS) portfolio, GCS can help reduce anthropogenic emissions from industrial and other hard-to-abate sources and support net-zero pathways alongside rapid emissions reductions[11, 25, 36]. Although large-scale GCS deployment remains at an early stage, projects are increasingly being implemented at the basin scale, where geological formations span large areas and support long-term operations and large-volume storage [27, 7, 52]. These large-scale deployments often involve multiple injection sites, complex monitoring and regulatory requirements, and diverse stakeholder participation. In this context, multi-stakeholder carbon capture and storage (CCS) initiatives—where operators, landowners, regulators, and investors interact within geologically connected sites—are becoming increasingly common.
Recent basin-scale GCS studies emphasize that large-scale GCS may be limited by pressure buildup and interference among neighboring projects, motivating coordinated pressure-management strategies among operators. For example, Bump and Hovorka [9] introduced “pressure space” as a basin-scale limiting commodity for CCS and showed that pressure-limited resource estimates can materially reduce practical storage expectations relative to pore-volume-only calculations. Jahediesfanjani et al. [28] quantify pressure-limited dynamic storage capacity across multiple U.S. basins and demonstrate how brine extraction can expand pressure-limited capacity. Plampin et al. [44] provide basin-scale dynamic resource estimates highlighting regional constraints such as pressure/injectivity limits. These studies reinforce the need for decision frameworks that explicitly account for pressure constraints and multi-project interference when evaluating future basin-scale deployment scenarios.
@@
In a multi-stakeholder CCS setting, each stakeholder pursues distinct interests and goals [6]. For example, operators typically manage one or more injection/extraction wells with objectives such as maximizing CO2 storage or financial returns. Landowners focus on maximizing land use value and royalty income, while regulators prioritize long-term safety, pressure containment, and environmental compliance, e.g. Class VI well regulations [48, 32]. Although only operators physically interact with the reservoir, interests of all other stakeholders also shape the operational landscape. For instance, regulators may impose safety constraints that limit injection volumes and rates, whereas landowners and operators may both seek to exploit the available pressure space [9] for maximum economic return. Moreover, the interconnected nature of subsurface dynamics means that one operator’s injection strategy can alter reservoir pressure and plume behavior, influencing the economic and operational outcomes of other operators.
@@
These inter-dependencies pose significant challenges to effective basin-wide coordination and optimization. Without a robust decision-making framework that integrates individual or collective objectives with system-wide regulatory constraints, stakeholders may risk adopting locally optimal strategies that undermine collective performance, leading to underutilized reservoir capacity, reduced economic returns, or even regulatory noncompliance. To this end, coordinated multi-agent optimization is essential to balance return, safety, and fairness in basin-scale GCS deployment.
1.2 Modeling Multi-agent Optimization in GCS
Basin-scale GCS modeling studies increasingly emphasize that pressure buildup and inter-project interference can become the binding system constraint long before CO2 plumes physically interact [17, 20, 3], especially when multiple large injection projects operate in proximity. For example, basin-scale numerical experiments in the Illinois Basin [5] show that active pressure management via brine production can materially reduce pressure buildup and shrink regulatory Areas of Review (AoR), highlighting how operational choices at one site can affect feasibility and monitoring requirements at others. More recent multi-well, multi-project simulations [53] explicitly track the co-evolution of pressure fronts and CO2 plumes under various operational variables, demonstrating that interference can tighten allowable injection envelopes over time and motivating coordinated planning at the basin scale.
@@ 1.2 Modeling Multi-agent Optimization in GCS
Despite extensive basin-scale modeling studies, optimal management of basin-scale geological carbon storage remains challenging, especially when multiple stakeholders with respective goals and interests are involved. On the one hand, high fidelity models that predict the pressure evolution and plume migration are needed to ensure meeting safety constraints. Optimization of this system requires multiple forward evaluations of these models, which becomes infeasible when the computational resources are limited. A remedy to this is employing machine learning-based proxy models to amortize the computation costs [49, 23]. On the other hand, the literature on multi-agent optimization for GCS or broader subsurface communities is scarce. Existing literature either only explored multi-objective optimization of one single stakeholder on various CO2 storage projects [55, 40, 34] or constrained co-optimization of one single stakeholder [60, 37, 47]. Recently, [42, 43] proposed a cooperative game model with multi-objective optimization that enables coordinated CO2 injection strategies among multiple agents in basin-scale carbon storage projects. This framework provides a general map of possible solution strategies for all possible coalition structures, but no mechanism for determining for imposing one of them rather than another. Consequently, unless there is a single centralized decision maker who assigns operational strategies to all agents, the outcome remains uncertain for all agents involved.
@@ 1.3 Constrained Markov Game (CMG) and Safe Multi-agent Reinforcement Learning (safe MARL)
In addition to the traditional optimization framework, which based on gradient-based or evolutionary algorithms, a framework referred to as constrained Markov game (CMG) [2] is suitable for constrained optimization and offers real-time optimized solutions when multiple actors are involved. A constrained Markov game extends the Markov game framework by adding safety constraints that all agents must comply with while optimizing their own objectives. In a CMG, the system is represented as a sequential decision-making process defined by states, actions, transitions, and rewards, along with explicit safety constraints that limit feasible actions. Safe multi-agent reinforcement learning provides practical methods to learn policies in such settings, ensuring agents maximize rewards while avoiding constraint violations. The detailed formulation of this approach is presented in section 3 and section 4.
@@ 1.4 Contributions and Paper Organization
This work develops a modeling and decision-support framework to analyze potential basin-scale, multi-operator GCS deployment scenarios and associated pressure-management challenges. The contributions of this study are twofold. First, we develop a safe multi-agent reinforcement learning framework tailored for multi-stakeholder, multi-site CO2 storage problems when each stakeholder has its own goals and interests. The framework accommodates these divergent interests while ensuring safe operations, and its performance is validated against traditional multi-objective optimization approaches. Second, a reward and penalty model is meticulously designed for CO2 storage projects, and various coalition structures based on shared rewards and penalties have been investigated to evaluate their impacts on stakeholder objectives and overall system performance.
The paper is organized as follows. Section 2 reviews the subsurface dynamics of GCS projects, including key objectives, operational constraints, and potential coalition strategies among stakeholders. Section 3 introduces fundamental concepts in reinforcement learning and Markov games, establishing the theoretical basis for our approach. Section 4 presents the proposed methodology and workflow tailored for GCS. Section 5 reports the case study results. Finally, the paper concludes with a discussion of key findings and a summary of contributions and directions for future work.
@@ 2.1 Overview of Basin-Scale GCS System
Geological CO2 storage typically consists of studies and analyses across various scales [22, 38] and involves complex coupled multi-physical processes [39, 29]. Deep saline aquifers have abundant storage potential and are regarded as reliable and feasible storage sites. When CO2 is compressed and supercritical CO2 is being injected into deep saline aquifers, a compositional multi-phase system is formed. In our study, we consider a compositional porous flow process where the governing equation of the system can be described as:
@@
After spatial and temporal discretizations, the above governing equations that describe the compositional flow in porous media can be reformulated in state-space form, which corresponds to the environment dynamics used in the subsequent Markov Decision Process (MDP) / Markov game formulations:
@@
where $t\in\{0,1,\dots,T-1\}$ is the discrete time index and $t+1$ denotes the next time step. $x_{t}\in\mathbb{R}^{2*N_{b}}$ is the system state (e.g., pressure and CO2 composition fields), $u_{t}\in\mathbb{R}^{N_{u}}$ is the control/action (e.g., well injection/extraction rates), and $y_{t}\in\mathbb{R}^{N_{obs}}$ denotes the simulated outputs/observations of interest.Here, $N_{b}$, $N_{u}$ and $N_{obs}$ denote the number of grid blocks, the dimension of the control variables, and the dimension of observation data, respectively. The mappings $f(\cdot)$ and $g(\cdot)$ represent the nonlinear state transition and observation operators induced by the discretized governing equations. Because the state dimension can be very large (tens of thousands to millions), using $x_{t}$ directly for reinforcement learning training is computationally prohibitive.
To address the computational burden of high-fidelity reservoir simulations, we leverage a surrogate reduced-order model based on the Embed-to-Control (E2C) architecture [51, 30, 16, 12, 15], which captures essential system dynamics while significantly reducing simulation time. Employing the machine learning-based reduced-order model, equation (3) can be approximated as:
@@
where $\mathrm{Enc}(\cdot;\varphi)$ is an encoder network that maps the high-dimensional reservoir state $x_{t}$ to a low-dimensional latent state $z_{t}\in\mathbb{R}^{n_{z}}$, and $\mathrm{Dec}(\cdot;\theta)$ is a decoder network that reconstructs the state $\hat{x}_{t}$ from $z_{t}$. The latent-space transition model $f_{z}(\cdot;\omega_{1})$ approximates the latent dynamics $z_{t+1}=f_{z}(z_{t},u_{t},\omega_{1})$, while the latent-space observation model $g_{z}(\cdot;\omega_{2})$ predicts the outputs $\hat{y}_{t+1}=g_{z}(z_{t+1},u_{t};,\omega_{2})$. The parameters $\varphi$, $\theta$, $\omega_{1}$, and $\omega_{2}$ denote the weights of the corresponding neural networks.
The E2C networks are trained offline using state–action trajectories generated from the high-fidelity reservoir simulator, with an illustrative example shown in Figure 12. Training minimizes a weighted objective of the following form:
$\mathcal{L}_{\mathrm{E2C}}=\mathcal{L}_{\mathrm{rec}}(x_{t},\hat{x}_{t})+\beta\,\mathcal{L}_{\mathrm{reg}}(z_{t},z_{t+1},u_{t})+\eta\,\mathcal{L}_{\mathrm{obs}}(y_{t+1},\hat{y}_{t+1}),$ (5)
where $\mathcal{L}_{\mathrm{rec}}$ is a reconstruction loss, $\mathcal{L}_{\mathrm{reg}}$ is a regularization term that promotes locally linear latent dynamics, and $\mathcal{L}_{\mathrm{obs}}$ penalizes mismatch between simulated and predicted observations. The scalars $\beta$ and $\eta$ are are weighting coefficients. Once trained, the surrogate replaces expensive simulator runs during MARL training by rolling out latent dynamics and decoding the required outputs. Prior studies [12, 13]have shown that this proxy model maintains high robustness while achieving desirable accuracy compared to high-fidelity simulations.
@@ 2.2 Stakeholders Objectives and Operational Conflicts
Basin-scale carbon capture and storage (CCS) projects involve a diverse array of stakeholders whose interests, responsibilities, and operational controls are inherently interdependent. Figure 1 illustrates the key actors and their roles in a typical CCS project. At the center of the system are operators, who manage CO2 injection wells (possibly also some brine extraction wells) and directly interact with the subsurface reservoir. Their primary objectives typically include maximizing CO2 storage volumes and financial returns through strategic control of injection rates and well operations.
@@
These stakeholders operate within a shared geologic basin and sometimes have misaligned goals. For instance, while both operators and landowners may favor maximizing pressure space for economic benefit, regulators may impose strict injection limits to prevent caprock fracturing or fluid migration beyond permitted boundaries. Moreover, one operator’s injection strategy can affect reservoir pressure distributions and CO2 plume migration patterns, potentially undermining neighboring operators’ performance or violating shared safety constraints.
@@ 2.3 Mathematical Formulation of Objectives
In this study, we consider the primary objective to be the maximization of financial returns. Each agent’s financial outcome is quantified using present value (PV). Accordingly, for operator $i$, the present value—or immediate reward—can be expressed as::
$\text{PV}_{t}^{i}(u)=\Bigg($ $\sum_{j=1}^{N_{\text{inj}}^{i}}(R_{\text{credit}}-R_{\text{op}})*q_{t,j}^{i}(u)$ (6)
@@
$\text{NPV}^{i}=\sum_{t=0}^{T}\gamma_{t}\cdot\text{PV}_{t}^{i}$ (7)
@@
$\text{p}^{i}_{grid}\leq\alpha^{i}*p^{i}_{frac}$ (8)
@@ 2.3.1 Multi-objective optimization formulation
Intuitively, optimal CO2 storage under multi-stakeholder settings can be formulated as a constrained multi-objective optimization (CMOO) problem, where the objective is to maximize all companies’ NPVs while adhering to the pressure constraints. The constrained multi-objective optimization problem can be formulated as:
$\max_{u}$ $\left\{\text{NPV}^{1}(u),\,\text{NPV}^{2}(u),\,\ldots,\,\text{NPV}^{n}(u)\right\}$ (9)
@@
$p_{grid}^{i}(u)\leq\alpha*p_{frac}^{i},i=1,2,.,n$
@@
where $NPV^{i}$ indicates the NPV of agent $i$, and the decisions $u$ are bounded by operational limits, pressure constraints, and the governing system transition dynamics.
@@
Figure 2: Sketch plot of MOO with two decision variable and two objectives. The gray areas represents the feasible decision and solution space. The blue dots in the objective space are Pareto optimal solutions forming the Pareto front, where improving one objective (an agent’s reward) necessarily degrades the other. Adapted from [41]
@@
Traditionally, CMOO problems are addressed using evolutionary algorithms to obtain a Pareto front (see Figure 2), which provides a spectrum of optimal trade-offs for decision-making. One widely used method is the Non-dominated Sorting Genetic Algorithm II (NSGA-II) [18], implemented using the PyMOO framework in this work [8]. In this approach, the algorithm iteratively evolves a population of solutions to converge toward the Pareto-optimal set. Each point on the Pareto front represents a trade-off where no objective can be improved without degrading another. This front can be used for post-optimization solution selection, typically choosing a “knee point” or preference-driven solution.
@@
Despite its effectiveness, the conventional CMOO framework may not fully capture the operational realities of CCS projects, as it assumes a single centralized decision maker with complete knowledge of basin-scale reservoir conditions, inter-well interactions, and the operational strategies of all stakeholders. In practice, CCS operations often involve multiple independent operators, each with limited information and potentially competing interests, making such centralized control unrealistic. Furthermore, CMOO typically produces a set of non-dominated solutions (a Pareto front). A separate decision rule is generally required to select a single implementable operating strategy.In multi-operator CCS settings, this introduces three practical limitations: (1). constructing a well-resolved Pareto front can be numerically expensive (and in some cases prohibitively so) because it requires many forward model evaluations. (2). the Pareto front enumerates all possible solutions under centralized optimization, but it does not specify what outcome will actually occur under decentralized decision making. (3). commonly used point-selection approaches (e.g., knee-point selection or preference-weighted scalarization) implicitly encode a choice criterion that may advantage particular stakeholders, and may not align with what is operationally feasible given contractual, regulatory, or organizational constraints These limitations reduce the representativeness of conventional CMOO for modeling real multi-operator CCS decision-making environments.
2.3.2 Constrained Markov game formulation